The line passing through the points (5, 1, a) and (3, b, 1) crosses the yz-plane at the
origin then find the a & b

Grade:upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
The line passing through the points (5, 1, a) and (3, b, 1) is
(x-5)/-2=(y-1)/b-1=(z-a)/1-a=k
So we get x=5-2k y=(b-1)k+1 z=(1-a)k+a
Also given that the line crosses plane at origin(0,0,0)
So we get k=5/2 by substituting x=0
a=5/3 and b=3/5 by substituting k=5/2 and y=0,z=0 respectively